c) You plan to make a series of deposits in an interest bearing account. You will



deposit R1000 today, R2000 in two years, and R8000 in five years. If you withdraw



R3000 in three years and R5000 in seven years, how much will you have after eight



years if the interest rate is 9%? (Assume annual compounding).

1
Expert's answer
2022-03-09T14:01:01-0500

Answer

Answer: 

Explanation: DEPOSIT NOW  $1000 * FVIF 9%,8 PERIODS = $1000 * 1.9926 = $1992.6

IN 2 YEARS = $2000 * FVIF 9%,6 PERIODS = $2000 * 1.6771 = $3354.20

IN 5 YEARS = $8000 * FVIF 9%, 3 PERIODS = $8000*1.2950 = $10360

WITHDRAWAL: IN 3 YEARS = ($3000) * FVIF 9%, 5 PERIODS = ($3000) * 1.5386 = ($4615.80)

IN 7 YEARS = ($5000) * FVIF 9%, 1 PERIOD = ($5000) * 1.0900 = ($5450)

Total value = $1992.6 + $3354.20 + $10360 - $4615.80 - $5450

Total value = $5,641

So, the total future value after eight years is $5,641